Melinda has a scrap piece of fabric with one angle measuring 52°. If she needs to find a complementary angle, what angle measure does she need to cut the next piece?
Answer by Theo(13342)

complementary angle to 52 = 90 - 52 = 38.

if two angles are complementary, their sum is 90 degrees.

52 + 38 = 90.